### v1

# Estimating effects of longitudinal modified treatment policies (LMTPs) on rates of change in health outcomes

Abstract
Longitudinal data often contains time-varying outcomes measured at multiple visits and scientific interest may lie in quantifying the effect of an intervention on an outcome’s rate of change. For example, one may wish to study the progression (or trajectory) of a disease over time under different hypothetical interventions. We extend the longitudinal modified treatment policy (LMTP) methodology introduced in D´ıaz et al. (2023) to estimate effects of complex interventions on rates of change in an outcome over time. We exploit the theoretical properties of a nonparametric efficient influence function (EIF)-based estimator to introduce a novel inference framework that can be used to construct simultaneous confidence intervals for a variety of causal effects of interest and to formally test relevant global and local hypotheses about rates of change. We illustrate the utility of our framework in investigating whether a longitudinal shift intervention affects an outcome’s counterfactual trajectory, as compared with no intervention. We present results from a simulation study to illustrate the performance of our inference framework in a longitudinal setting with time-varying confounding and a continuous exposure.
1 Introduction
Longitudinal data often contains time-varying outcomes measured at each visit. Many methods for causal inference from longitudinal data aim to estimate effects on an outcome at the end of the study, effectively ignoring any additional outcome information that is available or treating intermediate outcomes as time-varying confounders. However, there are settings where the scientific interest lies not in an end-of-study effect but rather in the effect of an intervention on the outcome’s rate of change. For example, one may wish to investigate the effect of an intervention on the rate of progression of a disease, but the methods available to answer such questions are mostly limited to parametric models (e.g., linear models for the rate of change). In this work, we build on the flexible and nonparametric longitudinal modified treatment policy (LMTP) methodology to answer causal research questions that target effects on rates of change in outcomes over time.
LMTPs quantify the effects of a broad class of interventions, including interventions that depend on the natural value of exposure (D´ıaz et al., 2023; Hoffman et al., 2024). The natural value refers to the value that a time-varying exposure would take at some time if the intervention was discontinued right before that time. LMTPs generalize “shift” interventions (e.g., where the natural value of exposure is shifted by a constant at every visit), “threshold” interventions (e.g., where the natural value of exposure is set to a threshold value when it fails to attain that threshold), and “stochastic” interventions (e.g., where the natural population- or individual-level values of exposure are modified) (Robins et al., 2004; Taubman et al., 2009; Muñoz and Van Der Laan, 2012; Haneuse and Rotnitzky, 2013; Young et al., 2014; D´ıaz and van der Laan, 2018; Kennedy, 2019; Sani et al., 2020). Static and dynamic regimes are also able to be studied as a special case. Since the LMTP framework accommodates such a wide range of interventions, it constitutes an ideal foundation to build upon. The LMTP approach has several other desirable qualities, particularly for longitudinal data. First, an LMTP approach allows researchers to evaluate policy-relevant interventions and encourages researchers to formulate estimands that correspond to interesting, complex research questions. Second, the LMTP methodology that is already developed properly accounts for time-varying confounding, a complication that is often encountered in longitudinal data. Third, the LMTP framework affords flexibility in the type of exposures being considered, allowing for continuous and categorical exposures as well as binary ones. Finally, the LMTP approach facilitates a focus on estimands that are more likely to satisfy the positivity assumption, which is key to reliable casual inference.
In this paper, we extend the LMTP estimation methodology to estimate effects on rates of change in an outcome over time and develop a comprehensive framework to conduct relevant inference. The paper is organized as follows: Section 2 provides an introduction to LMTPs; Section 3 formalizes causal effects on rates of change; Section 4 proposes an inference framework and outlines its usefulness in investigating whether the rate of change in an outcome changes when an LMTP is implemented; Section 5 investigates the performance of the framework in a longitudinal setting using simulated data; and finally, Section 6 discusses limitations of the framework and proposes some future areas of work.
2 LMTPs
In this section, we provide an introduction to LMTPs and establish relevant notation. Our setup is similar to D´ıaz et al. (2023) but extended to accommodate a continuous time-varying outcome. Consider a sample of independent and identically distributed (iid) observations $Z_{1},\ldots,Z_{n}\sim P$, where distribution $P$ is contained in a nonparametric statistical model. Let $Z=(L_{1},A_{1},Y_{1},\ldots L_{\tau},A_{\tau},Y_{\tau})$. For discrete time $t\in\{1,\ldots,\tau\}$, $L_{t}$ is a vector of time-varying covariates, $A_{t}$ is a time-varying exposure, and $Y_{t}$ is a continuous time-varying outcome. For a random variable $X$, denote its history and future as $\overline{X}_{t}=(X_{1},\ldots,X_{t})$ and $\underline{X}_{t}=(X_{t},\ldots,X_{\tau})$, respectively. For notational simplicity, we use $\overline{X}$ and $\underline{X}$ to denote the complete history and future. Denote the history of all random variables until just before $A_{t}$ as $H_{t}=(\overline{A}_{t-1},\overline{L}_{t},\overline{Y}_{t-1})$.
The causal model is formalized via a non-parametric structural equation model. This means data is generated based on deterministic functions $f_{L_{t}}$, $f_{A_{t}}$, and $f_{Y_{t}}$, such that
$L_{t}=f_{L_{t}}(A_{t-1},Y_{t-1},H_{t-1},U_{L,t})$
$A_{t}=f_{A_{t}}(H_{t},U_{A,t})$
$Y_{t}=f_{Y_{t}}(A_{t},H_{t},U_{Y,t}),$
where $U=(U_{L,t},U_{A,t},U_{Y,t}:t\in\{1,\dots,\tau\})$ is the set of exogenous variables. This implies a strict time-ordering at each time $t$ (i.e., $L_{t}\rightarrow A_{t}\rightarrow Y_{t}$).
We define an intervention as replacing $A_{t}$ in the structural model with $A_{t}^{d}$. Intervening on all exposures up until time t1t-1 is denoted by $\overline{A}^{d}_{t-1}=(A_{1}^{d},\ldots,A_{t-1}^{d})$ and induces the following counterfactual variables
$L_{t}(\overline{A}^{d}_{t-1})=f_{L_{t}}(A^{d}_{t-1},Y_{t-1}(\overline{A}^{d}_{t-1}),H_{t-1}(\overline{A}^{d}_{t-2}),U_{L,t})$
$A_{t}(\overline{A}^{d}_{t-1})=f_{A_{t}}(H_{t}(\overline{A}^{d}_{t-1}),U_{A,t})$
$Y_{t-1}(\overline{A}^{d}_{t-1})=f_{Y_{t}}(A^{d}_{t-1},H_{t}(\overline{A}^{d}_{t-1}),U_{Y,t}),$
where $H_{t}(\overline{A}^{d}_{t-1})=(\overline{A}^{d}_{t-1},\overline{L}_{t}(\overline{A}^{d}_{t-1}),\overline{Y}_{t-1}(\overline{A}^{d}_{t-1}))$ is the counterfactual history. We refer to $A_{t}(\overline{A}^{d}_{t-1})$ as the natural value of exposure; that is, the value that the time-varying exposure would take at time $t$ if the intervention was discontinued right before time $t$. The counterfactual outcome $Y_{t}(\overline{A}^{d}_{t})$ can be interpreted similarly.
Definition 1 (LMTP).
The intervention $A_{t}^{d}$ is called a LMTP if it is defined as $A_{t}^{d}=d(A_{t}(\overline{A}^{d}_{t-1}),H_{t}(\overline{A}^{d}_{t-1}))$ for a user-given function $d$.
An LMTP is defined by function $d$ that must take both the natural value of exposure and counterfactual history as inputs. One example of a common LMTP is an additive or multiplicative shift intervention, in which the natural value of exposure is modified by some constant $\delta$ at each time point. Rather than shifting all individuals, we may consider shifting individuals such as in Example 1 so that the shifted values fall inside the range of the empirical values, ensuring the positivity assumption holds by design.
Example 1 (Additive Shift LMTP).
Let $A_{t}$ be air pollution concentration at time $t$. Suppose $u_{t}$ exists such that $P(A_{t}>u_{t}|H_{t}=h_{t})=1$ for all $t\in\{1,\ldots,\tau\}$. For constant $\delta$, define
$d(a_{t},h_{t})=\begin{cases}a_{t}-\delta,&\text{ if }a_{t}-\delta\geq u_{t}(h_{t})\\ a_{t},&\text{ if }a_{t}-\delta<u_{t}(h_{t})\end{cases}.$
which can be interpreted as reducing the air pollution concentration input by $\delta$ only if the resulting shifted value falls above the minimum of the empirical values.
In this paper, we will often refer to this kind of shift intervention as our LMTP of interest; however, the theory developed applies to any LMTP that satisfies Definition 1. Other examples of LMTPs, including threshold interventions and LMTP stochastic interventions, are discussed in D´ıaz et al. (2023) and Hoffman et al. (2024).
3 Causal effects on rates of change
For a particular time $t$, consider the following causal estimand
$\theta_{t}=E[Y_{t}(\overline{A}^{d}_{t})],$
which is the expected counterfactual outcome at time $t$ under an intervention defined by function $d$. We use $\theta^{\prime}_{t}$ and $\theta^{\prime\prime}_{t}$ to distinguish estimands corresponding to two different interventions $d^{\prime}$ and $d^{\prime\prime}$. These functions can be any type of intervention, including no intervention.
Definition 2.
The vector of causal estimands across time $t\in\{1,\ldots,\tau\}$, denoted as $\overline{\theta}=(\theta_{1},\ldots,\theta_{\tau})$, is called the counterfactual outcome trajectory. If $d$ is no intervention, we call this vector the natural outcome trajectory to emphasize that it involves no counterfactuals.
For some intervention $d$, $\theta_{t}-\theta_{1}$ represents the change in the expected counterfactual outcome in the period from baseline to time $t$. Thus, we can define a relevant causal effect on the outcome’s rate of change, comparing two interventions $d^{\prime}$ and $d^{\prime\prime}$, as
$\Delta_{t}=\theta^{\prime\prime}_{t}-\theta^{\prime\prime}_{1}-(\theta^{\prime}_{t}-\theta^{\prime}_{1}).$
In other words, this is the difference in the change in the outcome, on average, in the period from baseline to time $t$, comparing $d^{\prime\prime}$ to $d^{\prime}$. For simplicity, we let $d^{\prime}$ be no intervention and $d^{\prime\prime}$ be a LMTP intervention for the rest of this paper, so our interest is how an LMTP changes the counterfactual outcome trajectory, as compared with no intervention.
Alternative definitions of $\Delta_{t}$ could also be considered. For example, $\Delta_{t}$ could be defined to compare adjacent time points by swapping out $\theta^{\prime\prime}_{1}$ and $\theta^{\prime}_{1}$ for $\theta^{\prime\prime}_{t-1}$ and $\theta^{\prime}_{t-1}$, respectively. The framework we propose in the next section will allow for flexibility in the way $\Delta_{t}$ is defined.
Definition 3.
The vector of causal effects for all combinations of time points $t\in\{2,\ldots,\tau\}$ with baseline time $t=1$ is $\Delta=(\Delta_{2},\ldots,\Delta_{\tau})$.
Theory for identification and nonparametric estimation of the vector of causal effects $\Delta$ follows straightforwardly from results established for any $\theta_{t}$ in D´ıaz et al. (2023). In brief, assumptions of positivity and strong sequential randomization are needed for identification based on a sequential regression formula. Two efficient influence function (EIF)-based estimators with many useful properties are proposed: the sequential doubly robust (SDR) estimator and the targeted minimum loss-based estimator (TMLE). Both of these estimators are $\sqrt{n}$-consistent and asymptotically normal. For these theoretical properties to hold, an additional assumption is needed: either $A_{t}$ must be discrete for all $t$ or if $A_{t}$ is continuous, $d(a_{t},h_{t})$ must be piecewise smooth invertible with respect to $a_{t}$. The framework that we propose in the next section relies heavily on the theoretical properties of these estimators. Nuisance functions in both the SDR estimator and TMLE are estimated using flexible data-adaptive techniques, such as Super Learner, that reduce the risk of model misspecification while still preserving proper convergence rates (Van der Laan et al., 2007). Additionally, these estimators are “doubly robust”, which refers to their ability to remain consistent even if some nuisance components are inconsistently estimated. Our implementation for estimating $\Delta$ follows straightforwardly from the functions available to estimate any $\theta_{t}$ in the lmtp R package (Williams and D´ıaz, 2023). In particular, we can estimate $\Delta$ by running the estimation algorithm for $\theta_{1},\dots,\theta_{\tau}$ (i.e., a loop of $\tau$ estimation runs) for both $d^{\prime}$ and $d^{\prime\prime}$.
4 Inference framework
So far, we have been able to rely on the existing LMTP infrastructure to estimate the vector of causal effects $\Delta$. However, additional theory and implementation is needed to conduct relevant inference for this vector. In this section, we propose a comprehensive inference framework that draws inspiration from Hothorn et al. (2008). We outline global and simultaneous inference that can be conducted with this framework that is relevant for investigating whether the rate of change in an outcome differs under an LMTP intervention compared with no intervention. Specifically, we use this framework to formally test for both global and local differences in the rate of change in the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ versus the natural outcome trajectory $\overline{\theta}^{\prime}$ and to construct simultaneous confidence intervals for the vector of causal effects $\Delta$, which quantifies the magnitude of difference.
4.1 Building intuition
Figure 1 provides intuition for how we construct our inference framework. Using a synthetic data example, we plot the corresponding natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ for comparison in Panel A. Our goal is to investigate whether these two trajectories differ in terms of their rate of change. In this example, there is a clear difference, as the two trajectories are non-parallel, with $\overline{\theta}^{\prime}$ declining more rapidly than $\overline{\theta}^{\prime\prime}$. First notice that when we take the difference between the two trajectories, we obtain the difference trajectory, as shown in Panel B, and find that this trajectory is non-constant over time. Subtracting each post-baseline point of the difference trajectory from the baseline point, we obtain the vector of causal effects $\Delta$, as shown in Panel C, and find that the components of $\Delta$ are non-zero. Therefore, one simple approach for globally testing whether two estimated trajectories differ significantly in terms of their rates of change would be to test whether at least one component of the estimated $\Delta$ vector is non-zero. If this global test is significant, we may proceed with local tests, individually testing which components of the estimated $\Delta$ are non-zero and providing these estimates and their simultaneous confidence intervals to quantify the magnitude of difference.
Figure 1: Synthetic data example comparing a potential natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$, in which a difference in the rates of change exists. Panel A visualizes the two decreasing trajectories. Panel B visualizes the difference trajectory obtained by subtracting $\overline{\theta}^{\prime}$ from $\overline{\theta}^{\prime\prime}$. Panel C visualizes the vector of causal effects $\Delta$ obtained by subtracting each point of the difference trajectory by the baseline point.
Since our target for inference, the vector of causal effects $\Delta$, is comprised of linear combinations of estimands, we formulate our inference framework in terms of linear functions of estimands using matrix algebra. We use this more general formulation because it affords the investigator flexibility in the way $\Delta$ is defined. Since many useful causal quantities are in fact just linear combinations of estimands, this makes our framework applicable for conducting inference for other causal quantities that may be of interest but are not the focus of this paper.
4.2 Set up
Consider stacking the natural outcome trajectory $\overline{\theta}^{\prime}$ and the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ into a single vector $\theta\in\mathbbm{R}^{2\tau\times 1}$ such that
$\theta=\begin{pmatrix}\theta^{\prime}_{1}&\cdots&\theta^{\prime}_{\tau}&\theta^{\prime\prime}_{1}&\cdots&\theta^{\prime\prime}_{\tau}\end{pmatrix}^{T}.$
Suppose we focus on constructing an estimator for $\theta$. For the purpose of illustrating this framework, we choose to use the SDR estimator instead of the TMLE because the SDR estimator is more robust to model misspecification in the sense that is is sequentially doubly robust, while the TMLE is only doubly robust (D´ıaz et al., 2023; Hoffman et al., 2024). Sequential double robustness guarantees that the estimator is consistent if either the outcome regression or treatment mechanism is consistently estimated for each time $t$. In contrast, double robustness only guarantees that the estimator is consistent if for some time $s$, all outcome regressions for $t>s$ and all treatment mechanisms for $t\leq s$ are consistently estimated. The only disadvantage of the SDR estimator is that unlike the TMLE, the SDR estimator can yield estimates of $\theta_{t}$ that fall outside the bounds of the parameter space. If this is an area of concern, one could consider using the TMLE instead. Even though the estimation algorithm for the TMLE differs from the SDR estimator, its construction based on the EIF allows a similar weak convergence result to be established so that the theory that we develop for inference still holds.
To construct a SDR estimator for any $\theta_{t}$, first define data transformation at time $s\in\{1,\ldots,t\}$
$\phi_{s,t}(z;\underline{\eta}_{s})=\sum_{p=s}^{t}\left(\prod_{k=s}^{p}r_{k}(a_{k},h_{k})\right)\left\{m_{p+1}(a_{p+1}^{d},h_{p+1})-m_{p}(a_{p},h_{p})\right\}+m_{s}(a_{s}^{d},h_{s}),$
which is dependent on the vector of nuisance parameters $\underline{\eta}_{s}=(r_{s},m_{s},\ldots,r_{t},m_{t})$, where $m$ denotes outcome regressions and $r$ denotes exposure density ratios. For $s=t,\ldots,1$, outcome regressions are defined recursively as
$m_{s}(a_{s},h_{s})=E\left[m_{s+1}(A_{s+1}^{d},H_{s+1})|A_{s}=a_{s},H_{s}=h_{s})\right],$
starting with $m_{t+1}(A_{t+1}^{d},H_{t+1})=Y_{t}$. Under the identification assumptions discussed earlier, $\theta_{t}=E\left[m_{1}(A_{1}^{d},L_{1})\right]$. For $s=1,\ldots,t$, the exposure density ratio is defined as
$r_{s}(a_{s},h_{s})=\frac{g_{s}^{d}(a_{s}|h_{s})}{g_{s}(a_{s}|h_{s})},$
where $g_{s}(a_{s}|h_{s})$ and $g_{s}^{d}(a_{s}|h_{s})$ are the densities of $A_{s}$ and $A_{s}^{d}$, respectively, conditional on $H_{s}=h_{s}$. Additional details on how the post-intervention density $g_{s}^{d}(a_{s}|h_{s})$ is defined can be found in D´ıaz et al. (2023).
The SDR estimator for any $\theta_{t}$ is constructed using a sequential regression algorithm (D´ıaz et al., 2023). This algorithm recursively regresses estimates of $\phi_{s+1,t}(Z;\underline{\eta}_{s})$ on $(A_{s},H_{s})$ for $s=t,\ldots,1$, starting with $\phi_{t+1,t}(Z;\underline{\eta}_{t+1})=Y_{t}$. The final output of this algorithm, $\phi_{1,t}(Z;\underline{\eta})$, is the (uncentered) EIF of $\theta_{t}$ in the nonparametric model. The SDR estimator for any $\theta_{t}$ under some intervention $d$ is then obtained by averaging the estimated EIF across all $i$ individuals, i.e.
$\widehat{\theta}_{t}=\frac{1}{n}\sum_{i=1}^{n}\phi_{1,t}(Z_{i},\underline{\widehat{\eta}}_{i}),$
In the special case where $d$ is no intervention, $\phi_{1,t}(Z_{i},\underline{\widehat{\eta}}_{i})=Y_{it}$, so the estimator is simply the sample mean of the outcome at time $t$. For notational simplicity, we use $\widehat{\phi}_{it}$ as shorthand for $\phi_{1,t}(Z_{i},\underline{\widehat{\eta}}_{i})$. The SDR estimator for $\theta$ is then given by $\widehat{\theta}_{n}\in\mathbbm{R}^{2\tau\times 1}$ such that
$\widehat{\theta}_{n}=\begin{pmatrix}\frac{1}{n}\sum_{i=1}^{n}\widehat{\phi}^{\prime}_{i1}&\cdots&\frac{1}{n}\sum_{i=1}^{n}\widehat{\phi}^{\prime}_{i\tau}&\frac{1}{n}\sum_{i=1}^{n}\widehat{\phi}^{\prime\prime}_{i1}&\cdots&\frac{1}{n}\sum_{i=1}^{n}\widehat{\phi}^{\prime\prime}_{i\tau}\end{pmatrix}^{T}.$
The SDR estimator $\widehat{\theta}_{n}$ satisfies a weak convergence result that follows directly from Theorem 4 in D´ıaz et al. (2023), which states that under proper convergence rates for the nuisance parameters, any $\widehat{\theta}_{t}$ weakly converges to a normal distribution with variance defined by the EIF.
Theorem 1 (Weak convergence of SDR estimator for $\theta$).
Assume conditions of Theorem 2 and 4 from D´ıaz et al. (2023) hold. Then,
$\sqrt{n}(\widehat{\theta}_{n}-\theta)\overset{d}{\longrightarrow}N_{2\tau}(0,\Sigma),$
where $0\in\mathbbm{R}^{2\tau\times 1}$ is a vector of zeros and $\Sigma\in\mathbbm{R}^{2\tau\times 2\tau}$ is a symmetric covariance matrix defined as
$\boldsymbol{\Sigma}=\begin{pmatrix}\sigma^{2}_{\phi^{\prime}_{1}}&&&\\ \sigma_{\phi^{\prime}_{2},\phi^{\prime}_{1}}&\sigma^{2}_{\phi^{\prime}_{2}}&&\\ \vdots&\vdots&\ddots&\\ \sigma_{\phi^{\prime}_{\tau},\phi^{\prime}_{1}}&\sigma_{\phi^{\prime}_{\tau},\phi^{\prime}_{2}}&\cdots&\sigma^{2}_{\phi^{\prime}_{\tau}}\\ \sigma_{\phi^{\prime\prime}_{1},\phi^{\prime}_{1}}&\sigma_{\phi^{\prime\prime}_{1},\phi^{\prime}_{2}}&\cdots&\sigma_{\phi^{\prime\prime}_{1},\phi^{\prime}_{\tau}}&\sigma^{2}_{\phi^{\prime\prime}_{1}}&\\ \sigma_{\phi^{\prime\prime}_{2},\phi^{\prime}_{1}}&\sigma_{\phi^{\prime\prime}_{2},\phi^{\prime}_{2}}&\cdots&\sigma_{\phi^{\prime\prime}_{2},\phi^{\prime}_{\tau}}&\sigma_{\phi^{\prime\prime}_{2},\phi^{\prime\prime}_{1}}&\sigma^{2}_{\phi^{\prime\prime}_{2}}&\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\\ \sigma_{\phi^{\prime\prime}_{\tau},\phi^{\prime}_{1}}&\sigma_{\phi^{\prime\prime}_{\tau},\phi^{\prime}_{2}}&\cdots&\sigma_{\phi^{\prime\prime}_{\tau},\phi^{\prime}_{\tau}}&\sigma_{\phi^{\prime\prime}_{\tau},\phi^{\prime\prime}_{1}}&\sigma_{\phi^{\prime\prime}_{\tau},\phi^{\prime\prime}_{2}}&\cdots&\sigma^{2}_{\phi^{\prime\prime}_{\tau}}\end{pmatrix}$,
where for some random variables $X$ and $W$, $\sigma^{2}_{X}=\text{Var}(X)$ and $\sigma_{X,W}=\text{Cov}(X,W)$.
Suppose $S_{n}\in\mathbbm{R}^{2\tau\times 2\tau}$ is a consistent estimator of $\frac{1}{n}\Sigma$, which means that
$nS_{n}\overset{p}{\longrightarrow}\Sigma.$ (1)
For example, $\Sigma$ may be estimated using the empirical covariance matrix of the estimated EIF matrix $\widehat{\phi}\in\mathbbm{R}^{n\times 2\tau}$, which is defined as
$\widehat{\phi}=\begin{pmatrix}\widehat{\phi}^{\prime}_{11}&\cdots&\widehat{\phi}^{\prime}_{1\tau}&\widehat{\phi}^{\prime\prime}_{11}&\cdots&\widehat{\phi}^{\prime\prime}_{1\tau}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ \widehat{\phi}^{\prime}_{n1}&\cdots&\widehat{\phi}^{\prime}_{n\tau}&\widehat{\phi}^{\prime\prime}_{n1}&\cdots&\widehat{\phi}^{\prime\prime}_{n\tau}\end{pmatrix}$.
Then, by Slutsky’s Theorem, the approximate distribution of $\widehat{\theta}_{n}$ is a multivariate normal distribution with mean $\theta$ and covariance matrix $S_{n}$, or
$\widehat{\theta}_{n}\overset{app}{\sim}N_{2\tau}(\theta,S_{n}).$ (2)
Now consider a linear function of $\theta$, denoted by
$\nu=K\theta\in\mathbbm{R}^{k\times 1},$
for some user-given constant matrix $K\in\mathbbm{R}^{k\times 2\tau}$.
Example 2 (Choosing $K$ to yield $\Delta$).
Let
[…]
Then,
$\nu=\begin{pmatrix}\theta^{\prime}_{1}-\theta^{\prime}_{2}-\theta^{\prime\prime}_{1}+\theta^{\prime\prime}_{2}\\ \theta^{\prime}_{1}-\theta^{\prime}_{3}-\theta^{\prime\prime}_{1}+\theta^{\prime\prime}_{3}\\ \vdots\\ \theta^{\prime}_{1}-\theta^{\prime}_{\tau}-\theta^{\prime\prime}_{1}+\theta^{\prime\prime}_{\tau}\end{pmatrix}=\begin{pmatrix}\theta^{\prime\prime}_{2}-\theta^{\prime\prime}_{1}-(\theta^{\prime}_{2}-\theta^{\prime}_{1})\\ \theta^{\prime\prime}_{3}-\theta^{\prime\prime}_{1}-(\theta^{\prime}_{3}-\theta^{\prime}_{1})\\ \vdots\\ \theta^{\prime\prime}_{\tau}-\theta^{\prime\prime}_{1}-(\theta^{\prime}_{\tau}-\theta^{\prime}_{1})\end{pmatrix}=\begin{pmatrix}\Delta_{2}\\ \Delta_{3}\\ \vdots\\ \Delta_{\tau}\end{pmatrix}$
which is the vector of casual effects $\Delta$.
Example 3 (Choosing $K$ to yield alternate version of $\Delta$).
Let
[…]
Then,
$\nu=\begin{pmatrix}\theta^{\prime\prime}_{2}-\theta^{\prime\prime}_{1}-(\theta^{\prime}_{2}-\theta^{\prime}_{1})\\ \theta^{\prime\prime}_{3}-\theta^{\prime\prime}_{2}-(\theta^{\prime}_{3}-\theta^{\prime}_{2})\\ \vdots\\ \theta^{\prime\prime}_{\tau}-\theta^{\prime\prime}_{\tau-1}-(\theta^{\prime}_{\tau}-\theta^{\prime}_{\tau-1})\end{pmatrix}$
which is an alternate definition of the vector of causal effects (comparing adjacent time points instead of comparing to baseline), that was introduced previously.
We can establish a weak convergence result for an estimator of $\nu$ that is analogous to Theorem 1. The SDR estimator for $\nu$ is given by $\widehat{\nu}_{n}=K\widehat{\theta}_{n}\in\mathbbm{R}^{k\times 1}$. By (2), the approximate distribution of $\widehat{\nu}_{n}$ is also a multivariate normal distribution. Specifically,
$\widehat{\nu}_{n}\overset{app}{\sim}N_{k}(\nu,S^{*}_{n}),$ (3)
where $S^{*}_{n}=KS_{n}K^{T}\in\mathbbm{R}^{k\times k}$ such that
$nS^{*}_{n}$ $=nKS_{n}K^{T}$
$=K(nS_{n})K^{T}$
$\overset{p}{\longrightarrow}K\Sigma K^{T}=\Sigma^{*}\in\mathbbm{R}^{k\times k}$ (4)
by (1). Consider the standardized quantity for $\widehat{\nu}_{n}$ given by
$T^{*}_{n}=(D^{*}_{n})^{-1/2}(\widehat{\nu}_{n}-\nu)\in\mathbbm{R}^{k\times 1}.$
An estimator for the correlation matrix of $T^{*}_{n}$ is given by $R^{*}_{n}\in\mathbbm{R}^{k\times k}$, which is defined as
$R^{*}_{n}=(D^{*}_{n})^{-1/2}S^{*}_{n}(D^{*}_{n})^{-1/2},$
where $D^{*}_{n}=\text{diag}(S^{*}_{n})\in\mathbbm{R}^{k\times k}$ is the diagonal matrix of $S^{*}_{n}$ such that
$nD^{*}_{n}$ $=n\text{diag}(S^{*}_{n})$
$=\text{diag}(nS^{*}_{n})$
$\overset{p}{\longrightarrow}\text{diag}(\Sigma^{*})=D^{*}\in\mathbbm{R}^{k\times k}$ (5)
by (4). Then, by (3), the approximate distribution of $T^{*}_{n}$ is once again a multivariate normal distribution. Specifically,
$T^{*}_{n}\overset{app}{\sim}N_{k}(0,R^{*}_{n}).$ (6)
Finally, notice that Slutsky’s Theorem can be used again since
$R^{*}_{n}$ $=(D^{*}_{n})^{-1/2}S^{*}_{n}(D^{*}_{n})^{-1/2}$
$=(nD^{*}_{n})^{-1/2}(nS^{*}_{n})(nD^{*}_{n})^{-1/2}$
$\overset{p}{\longrightarrow}(D^{*})^{-1/2}\Sigma^{*}(D^{*})^{-1/2}=R^{*}\in\mathbbm{R}^{k\times k}$
by (4) and (5) and that
$T^{*}_{n}$ $=(D^{*}_{n})^{-1/2}(\widehat{\nu}_{n}-\nu)$
$=(nD^{*}_{n})^{-1/2}\sqrt{n}(\widehat{\nu}_{n}-\nu)$
Theorem 2 (Weak convergence of SDR estimator for $\nu$).
Assume conditions of Theorem 1 hold. Then,
$T^{*}_{n}=(nD^{*}_{n})^{-1/2}\sqrt{n}(\widehat{\nu}_{n}-\nu)\overset{d}{\longrightarrow}N_{k}(0,R^{*}).$
4.3 Global inference
As previously described, we are interested in testing for a global difference in the rate of change of the natural outcome trajectory $\overline{\theta}^{\prime}$ versus the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$. We use Theorem 2 to establish a general global hypothesis testing procedure that can be used to conduct not only this particular global test but also other relevant global tests.
Consider the following global hypothesis
$H_{0}:\nu=h$
$H_{1}:\nu\neq h,$ (7)
where
$\nu=\begin{pmatrix}\nu_{1}&\nu_{2}&\cdots&\nu_{k}\end{pmatrix}^{T}\in\mathbbm{R}^{k\times 1}$
and
$h=\begin{pmatrix}h_{1}&h_{2}&\cdots&h_{k}\end{pmatrix}^{T}\in\mathbbm{R}^{k\times 1}$
is a user-given vector of constants that is oftentimes set equal to the zero vector 0.
Example 4 (Global test for difference in rate of change).
Let $K$ be the matrix in Example 2 and $h$ be the 0 vector. Then, the global null hypothesis becomes
$H_{0}:\Delta_{2}=\cdots=\Delta_{\tau}=0$
Testing this global hypothesis is equivalent to testing whether there is a global difference in the rate of change of the natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$.
To establish global tests for the hypothesis in (7), notice that under $H_{0}$,
$T^{*}_{n}=(D^{*}_{n})^{-1/2}(\widehat{\nu}_{n}-h)\overset{d}{\longrightarrow}N_{k}(0,R^{*})$ (8)
based on Theorem 2. Therefore, standard global hypothesis tests can be constructed. One such test that we consider is a Wald test. A Wald test follows trivially from (8) by application of Slutsky’s Theorem.
Corollary 1 (Global Wald test).
Suppose $R^{*}_{n}$ is a consistent estimator of the correlation matrix $R^{*}$. A Wald test statistic is defined as
$T^{*}_{w}=(T^{*}_{n})^{T}(R^{*}_{n})^{-1}(T^{*}_{n}).$
Under $H_{0}$ in (7), the limiting distribution of $T^{*}_{w}$ is a chi-square distribution with $k$ degrees of freedom $\chi^{2}_{k}$, or
$T^{*}_{w}\overset{d}{\longrightarrow}\chi^{2}_{k}.$
Let $t^{*}_{w}$ be the observed Wald test statistic. $H_{0}$ is rejected at the significance level $\alpha$ if $t^{*}_{w}>q_{w,\alpha}$, where $q_{w,\alpha}$ is the $(1-\alpha)$ quantile of $\chi^{2}_{k}$, i.e. $P(\chi^{2}_{k}\leq q_{w,\alpha})=1-\alpha$. The global p-value is calculated as $p_{w}=P(\chi^{2}_{k}>t^{*}_{w})$.
An alternative global test for the hypothesis in (7) that we also consider is a maximum test. Under $H_{0}$, consider
$T^{*}_{n}=\begin{pmatrix}T^{*}_{1,n}&T^{*}_{2,n}&\cdots&T^{*}_{k,n}\end{pmatrix}$,
where $T^{*}_{j,n}$ is the $j$th component of $T^{*}_{n}$. Denote the maximum of $|T^{*}_{j,n}|,j=1,\ldots,k$ by $\text{max}(|T^{*}_{n}|)$. Finally, notice that
$P(\text{max}(|T^{*}_{n}|)\leq t)$ $=P(|T^{*}_{n}|\leq t)$
$=P(-t\leq T^{*}_{n}\leq t).$
Therefore, the limiting distribution of $\text{max}(|T^{*}_{n}|)$ follows trivially from (8).
Corollary 2 (Global maximum test).
A maximum test statistic is defined as
$T^{*}_{m}=\text{max}(|T^{*}_{n}|)$
Under $H_{0}$ in (7), the limiting distribution of $T^{*}_{m}$ is given by
$P(T^{*}_{m}\leq t)\overset{d}{\longrightarrow}\int_{-t}^{t}\cdots\int_{-t}^{t}f(x_{1},\ldots,x_{k};R^{*}_{n})dx_{1}\cdots dx_{k}\coloneq g(t;R^{*}_{n})$
where $f(x_{1},\ldots,x_{k};R^{*}_{n})$ is the density for a multivariate normal distribution with mean 0 and correlation matrix $R^{*}_{n}$.
Let $t^{*}_{m}$ be the observed maximum test statistic. $H_{0}$ is rejected at the significance level $\alpha$ if $t^{*}_{m}>q_{m,\alpha}$, where $q_{m,\alpha}$ is the $(1-\alpha)$ quantile of the limiting distribution of $T^{*}_{m}$, i.e. $g(q_{m,\alpha};R^{*}_{n})=1-\alpha$. The global p-value is calculated as $p_{m}=1-g(t^{*}_{m};R^{*}_{n})$.
Although both the Wald and maximum tests are valid hypothesis tests to use for global inference, the latter test can be extended to construct simultaneous inference procedures that control the overall type I error without being overly conservative. We discuss this extension for simultaneous inference more in the next section. Although we choose to focus on a Wald test and maximum test, other global hypothesis tests can be developed within our framework.
4.4 Simultaneous inference
When the global null hypothesis is rejected, it is natural to proceed with testing each individual hypothesis to see where exactly the difference lies. For example, if we find a global difference in the rate of change comparing the natural outcome trajectory $\overline{\theta}^{\prime}$ with the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$, then we would want to test for local differences. We use results from Corollary 2 to establish a general local hypothesis testing procedure. Additionally, we extend these results to construct simultaneous confidence intervals.
Consider $k$ local hypotheses for $j=1,\ldots,k$
$H_{0}^{j}:\nu_{j}=h_{j}$
$H_{1}^{j}:\nu_{j}\neq h_{j},$ (9)
where $\nu_{j}$ and $h_{j}$ are the $j$th components of $\nu$ and $h$, respectively.
Consider $T^{*}_{j,n},j=1,\ldots,k$. Based on (8) and under $H^{j}_{0}$, a local test is given by
$T^{*}_{j,n}=(D^{*}_{n})_{jj}^{-1/2}(\widehat{\nu}_{j,n}-h_{j})\overset{d}{\longrightarrow}N(0,1),$ (10)
where $(D^{*}_{n})_{jj}$ is the matrix component corresponding the $j$th row and $j$th column of $D^{*}_{n}$ and is equal to $(S^{*}_{n})_{jj}$. This local test has rejection threshold $z_{\alpha/2}$, which is the $(1-\alpha/2)$ quantile of a standard normal distribution $Z$, i.e. $P(Z\leq z_{\alpha/2})=1-\alpha/2$. For $j=1,\ldots,k$, it yields observed test statistic $t^{*}_{j}$ with p-value $p_{j}=P(Z>|t^{*}_{j}|)$.
However, when these local tests are conducted simultaneously, the probability of falsely rejecting at least one true null hypothesis becomes larger than the nominal significance level $\alpha$. This is the origin of the multiple comparisons problem. To ensure that the overall type I error, also known as the family wise error rate, is bounded by $\alpha$, we can use a single-step multiple testing procedure to identify a new threshold for rejection and calculate adjusted p-values. Bounding the overall type I error by $\alpha$ means choosing a new rejection threshold $c_{\alpha}$ such that
$P\left(\left|t^{*}_{j}\right|\leq c_{\alpha}\ \text{for all}\ j\ \big{|}H_{0}\right)\geq 1-\alpha.$
We first consider a simple procedure that is frequently employed known as the Bonferroni procedure. To motivate the Bonferroni procedure, notice that the following statement
$1-\alpha=P\left(\left|t^{*}_{j}\right|\leq c_{b,\alpha}\ \text{for all}\ j\right)\geq 1-\sum_{j=1}^{k}P\left(\left|t^{*}_{j}\right|\leq c_{b,\alpha}\right).$
holds when $c_{b,\alpha}=z_{\alpha/(2k)}$. Thus, the Bonferroni-adjusted p-values are calculated by multiplying the unadjusted p-values with the number of comparisons $k$, i.e. $p_{b,j}=k\times p_{j},j=1,\ldots,k$.
Since the Bonferroni procedure is free of distributional assumptions and does not assume any dependence between tests, it is generally a conservative procedure, meaning that it oftentimes yields an overall type I error that is less than $\alpha$. This motivates us to consider an alternative multiple testing procedure that is generally less conservative. We refer to this alternative procedure as the maximum procedure since it is developed using results from Corollary 2. To motivate the maximum procedure, notice that
$1-\alpha$ $=P\left(\left|t^{*}_{j}\right|\leq c_{m,\alpha}\ \text{for all}\ j\right)$
$=P(\text{max}(|t^{*}_{1}|,\ldots,|t^{*}_{k}|)\leq c_{m,\alpha})$
$=P(t^{*}_{m}\leq c_{m,\alpha})$
$\approx g(c_{m,\alpha};R^{*}_{n}),$
with the approximation in the final line holding when $c_{m,\alpha}=q_{m,\alpha}$ by definition of the limiting distribution of $T^{*}_{m}$ in Corollary 2. Thus, the maximum-adjusted p-values are calculated as $p_{m,j}=1-g(|t^{*}_{j}|;R^{*}_{n}),j=1,\ldots,k$. We summarize both proposed single-step multiple testing procedures in Corollary 3.
Corollary 3 (Local test).
For $j=1,\ldots,k$, a local test statistic is defined as
$T^{*}_{j,n}.$
Under $H_{0}^{j}$ in (9), the limiting distribution of $T^{*}_{j,n}$ is a standard normal distribution $Z$, or
$T^{*}_{j,n}\overset{d}{\longrightarrow}Z.$
Let $t^{*}_{j},j=1,\ldots,k$ be the observed test statistics. Using the Bonferroni procedure, $H_{0}^{j}$ is rejected at the significance level $\alpha$ if $|t^{*}_{j}|>z_{\alpha/(2k)}$, and the p-values are calculated as $p_{b,j}=k\times P(Z>|t^{*}_{j}|),j=1,\ldots,k$. Using the maximum procedure, $H_{0}^{j}$ is rejected if $|t^{*}_{j}|>q_{m,\alpha}$, and the p-values are calculated as $p_{m,j}=1-g(|t^{*}_{j}|;R^{*}_{n}),j=1,\ldots,k$.
We can extend the results in Corollary 3 to construct simultaneous confidence intervals for $\nu_{j},j=1,\ldots,k$. We first consider the construction of pointwise confidence intervals, which can be done by inverting the local test for $H^{j}_{0}$ given in (10). Specifically, we invert
$P(|T^{*}_{j,n}|\leq z_{\alpha/2})=1-\alpha,$
where $T^{*}_{j,n}=(D^{*}_{n})_{jj}^{-1/2}(\widehat{\nu}_{j,n}-\nu_{j})$, and obtain the following pointwise confidence interval for $\nu_{j}$
$\bigg{(}\widehat{\nu}_{j,n}-z_{\alpha/2}\sqrt{(D^{*}_{n})_{jj}},\ \widehat{\nu}_{j,n}+z_{\alpha/2}\sqrt{(D^{*}_{n})_{jj}}\bigg{)}.$
A simple approach for constructing simultaneous confidence intervals for $\nu_{j},j=1,\ldots,k$ would be to replace the original threshold $z_{\alpha/2}$ with a larger threshold $c_{\alpha}$, i.e.
$\bigg{(}\widehat{\nu}_{j,n}-c_{\alpha}\sqrt{(S^{*}_{n})_{jj}},\ \widehat{\nu}_{j,n}+c_{\alpha}\sqrt{(S^{*}_{n})_{jj}}\bigg{)}.$
Both the Bonferroni and maximum rejection thresholds given in Corollary 3 would be valid choices for $c_{\alpha}$.
In this paper, we only consider two possible single-step procedures. Each of these procedures use a common threshold value for conducting simultaneous inference. Other single-step procedures, as well as step-wise procedures, which tend to be more powerful, may also be of interest.
5 Simulation study
We present a simulation study to illustrate the performance of our inference framework on balanced longitudinal data, where all individuals are assessed at the same set of pre-specified assessment times $v_{t}$, $t\in{1,\ldots,\tau}$. We considered a $\tau=4$ setting with time-varying confounding and a continuous exposure. We used the directed acyclic graph (DAG) in Figure 2 to generate 1000 datasets of each sample size $n\in\{250,500,1000,2500,5000,10000\}$ containing 3 continuous, time-varying variables: confounder $L$, exposure $A$, and outcome $Y$.
$L_{1}$$L_{2}$$L_{3}$$L_{4}$$A_{1}$$A_{2}$$A_{3}$$A_{4}$$Y_{1}$$Y_{2}$$Y_{3}$$Y_{4}$
Figure 2: DAG representing data generating process when $\tau=4$.
We used a data generating mechanism that yields non-linear, decreasing counterfactual and natural outcome trajectories. These trajectories decrease gradually at first and then begin to decrease more rapidly, as visualized in Figure 3, reflecting a trend commonly seen in health outcomes, such as cognitive decline. For illustration purposes, we considered a simple LMTP, in which the exposure input is shifted down by 1 unit at each $t$ for all individuals, i.e. $d(a_{t},h_{t})=a_{t}-1$. The data generating mechanism is as follows
$L_{1}\sim\texttt{Normal}\Big{\{}1,1\Big{\}}$
$A_{1}|L_{1}\sim\texttt{Normal}\Big{\{}8.5-L_{1},1\Big{\}}$
$Y_{1}|(L_{1},A_{1})\sim\texttt{Normal}\Big{\{}70.5+\gamma_{t}(-L_{1}+\alpha A_{1}),1\Big{\}}$
$L_{t}|(L_{t-1},A_{t-1},Y_{t-1})\sim\texttt{Normal}\Big{\{}5+0.47L_{t-1}-0.24A_{t-1}-0.05Y_{t-1}-0.3v_{t},1\Big{\}}\ \text{for}\ t\in\{2,3,4\}$
$A_{t}|(L_{t},A_{t-1},Y_{t-1})\sim\texttt{Normal}\Big{\{}10-0.2L_{t}+0.1A_{t-1}-0.05Y_{t-1}+0.5v_{t},1\Big{\}}\ \text{for}\ t\in\{2,3,4\}$
$\begin{split}Y_{t}|(L_{t},A_{t},Y_{t-1})\sim\texttt{Normal}\Big{\{}78+\gamma_{t}(-0.5L_{t}+\alpha A_{t}-0.15Y_{t-1})-0.3v_{t}-0.2v_{t}^{2}-0.1v_{t}^{3}\\ -\beta(0.1A_{t}v_{t}+0.04A_{t}v_{t}^{2}+0.02A_{t}v_{t}^{3}),1\Big{\}}\ \text{for}\ t\in\{2,3,4\}\end{split}$
where $\texttt{Normal}\{\mu,\sigma^{2}\}$ is a normal distribution with mean $\mu$ and variance $\sigma^{2}$. Here, $\gamma_{t}$ parametrizes outcome $Y$ so that the vector of causal effects $\Delta$ is exactly 0 when $\beta=0$, creating a setting where the natural outcome trajectory $\overline{\theta}^{\prime}$ and the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ are parallel and thus do no differ in terms of the rate of change. We calculate $\gamma_{t}$ for each $t\in\{1,\ldots,\tau\}$ recursively, starting with $\gamma_{1}$. More specifically, we analytically calculate the expected natural outcome $\theta_{1}^{\prime}$ and the expected counterfactual outcome $\theta_{1}^{\prime\prime}$ as if $\gamma=1$, calculate $\gamma_{1}=\frac{-\alpha}{\theta_{1}^{\prime\prime}-\theta_{1}^{\prime}}$, update $\theta_{1}^{\prime}$ and $\theta_{1}^{\prime\prime}$ based on this new $\gamma_{1}$, and repeat this entire process for the following $t\in\{2,\ldots,\tau\}$. Therefore, $\alpha$ represents the difference between $\theta_{t}^{\prime}$ and $\theta_{t}^{\prime\prime}$ at each $t$ in the setting where $\overline{\theta}^{\prime}$ and $\overline{\theta}^{\prime\prime}$ are parallel and there is no difference in the rate of change. Additionally, $\beta$ captures the magnitude of interaction between assessment time $v$ and exposure $A$ and indirectly controls how different $\overline{\theta}^{\prime}$ is from $\overline{\theta}^{\prime\prime}$ in terms of the rate of change. When $\beta=0$, the null hypothesis $H_{0}$ in (7) holds, as $\overline{\theta}^{\prime}$ and $\overline{\theta}^{\prime\prime}$ are parallel. When $\beta>0$, the alternative hypothesis $H_{1}$ holds, as $\overline{\theta}^{\prime}$ and $\overline{\theta}^{\prime\prime}$ are non-parallel, with higher values yielding greater differences in the rates of change. Therefore, $\beta$ can be thought of as an arbitrary “effect size” for the difference in the rate of change. Since all variables are generated using normal distributions and are associated linearly with each other, analytical calculation of the true $\overline{\theta}^{\prime}$ and $\overline{\theta}^{\prime\prime}$ follows straightforwardly by application of linearity of expectation. For our simulation, we let $\alpha=-2$ and $v_{t}\in\{0,2,4,6\}$.
We assessed the performance of our inference framework for investigating whether the natural outcome trajectory $\overline{\theta}^{\prime}$ differs from the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$ in terms of the rate of change. For each $n$, we repeated our simulation for a sequence of $\beta\in[0,1]$ to assess performance under both the null and alternative hypotheses. Figure 3 presents estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and the counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$, as well as the corresponding vector of causal effects $\Delta$, from a single simulated dataset when $\beta=1$ and $n=2500$. The simultaneous 95% confidence intervals of $\Delta$ capture the true non-zero causal effects, showing that our approach for estimation and simultaneous inference is performing as expected. Figure 4 presents the bias in the SDR estimator for $\Delta$ across several sample sizes $n$ when $\beta\in\{0,0.5,1\}$. As $n$ increases, bias converges to 0 as expected for each component of the SDR estimator of $\Delta$. Convergence is more rapid when $\beta$ is low and for components of the SDR estimator of $\Delta$ that are associated with earlier time points, suggesting that a larger sample size is needed for estimation using the Super Learner when there are stronger interaction effects and thus more variation in $Y_{t}$ in the underlying data generating process and when there are more time points to recurse through in the sequential regression algorithm used for estimation.
Figure 5 presents the power of the global Wald and maximum tests across a sequence of $\beta$ and for several $n$. Here, power is defined as the probability of rejecting the null hypothesis $H_{0}$. The result at $\beta=0$ is interpreted as the type I error since it indicates the probability that the null hypothesis is falsely rejected. Figure 6 presents results under no multiple testing correction and under the Bonferroni and maximum procedures for the simultaneous power of the local tests and the simultaneous coverage of the confidence intervals of the vector of causal effects $\Delta$. Here, simultaneous power is defined as the probability that $H_{0}^{j}$ is rejected for all $j=2,\ldots,\tau$, with the result at $\beta=0$ interpreted as the simultaneous type I error. Simultaneous coverage is defined as the probability that the confidence interval for $\Delta_{j}$ covers the true value for all $j=2,\ldots,\tau$. Subtracting the simultaneous coverage result at $\beta=0$ from 1 yields the overall type I error, or family wise error rate, as it corresponds to the probability of falsely rejecting $H_{0}^{j}$ for at least one local test. The results in Figures 5 and 6 highlight several important characteristics. First, as $\beta$ and $n$ increase, power increases as expected for both the global and local tests. Second, the global Wald test tends to be slightly more powerful than the global maximum test. Third, when the data generating mechanism yields local tests that are poorly correlated, such as in this simulation, then the maximum procedure displays no advantage over the Bonferroni procedure in terms of simultaneous power and coverage. Fourth, using the empirical covariance matrix of the estimated EIFs to conduct inference yields anti-conservative confidence intervals and global tests with slightly below nominal simultaneous coverage and above nominal type I error. Previous work has shown similar results in a longitudinal setting when using the empirical variance of the EIF as an estimator for the variance of doubly robust estimators like the SDR estimator and has highlighted that this behavior is exacerbated as positivity violations increase (Tran et al., 2023). This is because the covariance matrix and the variance are themselves estimands that have a non-negligible first order bias term, but plug-in estimators, such as the empirical covariance matrix, do no correct for this bias term.
Figure 3: Results from a single simulated dataset when $\beta=1$ and $n=2500$. Panel A visualizes estimates of the natural outcome trajectory $\overline{\theta}^{\prime}$ and counterfactual outcome trajectory $\overline{\theta}^{\prime\prime}$. Panel B visualizes an estimate of the difference trajectory (solid, black line) and the true difference trajectory (dashed, red line). Panel C visualizes an estimate of the vector of causal effects $\Delta$ and its simultaneous 95% confidence intervals (black), as well as the true vector of causal effects (red).
Figure 4: Bias of the SDR estimator for the vector of causal effects $\Delta$ across 1000 simulated datasets for a sequence of $n$ and $\beta\in\{0,0.5,1\}$.
Figure 5: Power of the global Wald (red) and maximum (blue) tests across 1000 simulated datasets for a sequence of $\beta$ and several $n$. The dashed, black line represents the threshold for a power of 0.80.
Figure 6: Simultaneous power of the local tests (Panel A) and simultaneous coverage of the confidence intervals of the vector of causal effects $\Delta$ (Panel B) under no multiple testing correction (green) and under the Bonferroni (red) and maximum (blue) procedures across 1000 simulated datasets for a sequence of $\beta$ and several $n$. The dashed, black line represents the threshold for a power of 0.80 (Panel A) or a coverage of 0.95 (Panel B).
6 Discussion
When analyzing longitudinal data, scientific interest may lie in investigating the effect of an intervention on an outcome’s rate of change. In a health context, this could entail studying the progression (or trajectory) of a disease over time under different hypothetical interventions. Here, we extend the LMTP methodology to estimate effects of complex interventions on rates of change in an outcome over time. We also propose a novel inference framework to formally test hypotheses about whether these interventions affect the outcome’s counterfactual trajectory. By building off the LMTP methodology, our approach is flexible and inherits many desirable properties already discussed. Our inference framework exploits the theoretical properties of the EIF-based SDR estimator to conduct global and local hypothesis tests and construct simultaneous confidence intervals for linear functions of $\theta$. Flexibility in the choice of the $K$ matrix enables our framework to be used for a variety of causal quantities and hypothesis tests of interest. We have illustrated the utility of our framework in investigating whether a longitudinal shift intervention changes the counterfactual outcome trajectory, as compared with no intervention.
Our approach has some limitations, and questions remain open for future work. First, the LMTP estimation procedure is computationally intensive. This is in part due to the use of the Super Learner for nuisance parameter estimation and could be mitigated by using a more computationally efficient data-adaptive technique. However, more fundamentally, it is due to the use of an estimation algorithm that recurses backwards through time and must be run $\tau$ times to estimate the entire counterfactual outcome trajectory under a specific intervention. Future work could consider constructing an alternative estimator that would allow estimation of the entire counterfactual outcome trajectory using a single algorithmic run. Additionally, variance estimation of EIF-based doubly robust estimators is conventionally achieved by taking the empirical variance of the EIF. Following this practice, we use the empirical covariance matrix of the EIFs as a plug-in estimator of the covariance matrix of the SDR estimator for $\theta$ in our inference framework. However, this plug-in estimator can yield anti-conservative confidence interval coverage and type I error, as shown in our simulation. Tran et al. (2023) demonstrated similar results when using the empirical variance of the EIF in the context of longitudinal dynamic treatment regimes and proposed two novel approaches for variance estimation. Future work could consider adapting one of these approaches to achieve more robust estimation of the covariance matrix in our inference framework. Finally, the LMTP methodology does not easily accommodate certain data complications that are commonly seen in longitudinal studies. One such complication is non-monotone missingness, which can occur when there are complex patterns of drop-out and re-entry. Although the LMTP methodology was extended to accommodate monotone loss-to-follow-up, where missingness at time $s$ implies missingness at all $t>s$, a similar extension does not exist for non-monotone missingness. Another complication is irregular assessment times, where the timing of the assessment varies among individuals from some pre-specified time. The current LMTP methodology requires coarsening the time scale into discrete intervals and it is unclear whether this coarsening approach always yields sensible results. In an follow-up paper, we will propose approaches to handling both non-monotone missingness and irregular assessment times when estimating LMTP effects.